# Dynamic Parametric Retrieval Augmented Generation for Test-time Knowledge Enhancement

### v2

@@ 1 Introduction
To mitigate incorrect internal knowledge and improve performance in QA, researchers have developed knowledge-augmented methods for LLMs. One of the most used methods called Retrieval-Augmented Generation (RAG) [13], which retrieves relevant documents from external sources (e.g., Wikipedia) and integrates both the retrieved documents and the query into the model’s input [17] (as shown in Figure 1 (a)). Standard RAG methods [3; 18; 37] share a common approach: injecting external knowledge into LLMs by appending retrieved documents to the input context, referred to as in-context injection. This technique harnesses the reasoning capabilities of LLMs to address knowledge gaps and reduce misinformation [5]. However, it introduces several challenges. As the number and length of retrieved documents increase, the extended context leads to significantly higher inference costs. Additionally, when the internal knowledge of LLMs conflicts with the external context, a phenomenon known as RAG hallucination (or knowledge conflict) arises, causing LLMs to produce incorrect answers even when accurate documents are provided [51; 40].
@@ 2.1 Retrieval Augmented Generation
Large language models (LLMs) have demonstrated remarkable performance across diverse applications. However, their inherent knowledge often falls short in handling knowledge-intensive tasks, highlighting the need for external knowledge integration to ensure robust performance in such contexts. A prominent approach to bridging this gap is Retrieval-Augmented Generation (RAG), which enhances LLMs by incorporating relevant external knowledge sources [4; 44; 45; 14]. The retrieved documents are appended to the LLM’s input context, enabling it to leverage knowledge beyond its training data [24]. We refer to this as the in-context injection paradigm. However, this approach leads to high inference costs and can exceed context length limits as the number and length of retrieved documents increase [48]. To address this issue, recent study introduces Parametric RAG (PRAG) [39], a paradigm that fine-tunes the model on augmented documents, encoding useful information into its parameters. While PRAG mitigates the inference cost, it introduces additional training and storage costs due to the need to obtain and store LoRA parameters. Our proposed method significantly reduces the high costs associated with standard RAG and PRAG while achieving superior generalization. By combining contextual knowledge with test-time generated parametric knowledge through DyPRAG, it enables LLMs to better manipulate knowledge, effectively mitigating RAG hallucination. This makes it a more practical and robust solution for real-world applications.
@@ 2.2 Context Compression
Context compression is widely adopted to improve the efficiency of LLMs in processing contextual knowledge. Recent studies [32; 11; 8] propose condensing long contexts into soft prompts, allowing LLMs to utilize information more effectively. Meanwhile, other works  [29; 46] focus on transforming context chunk into LoRA modules to improve the understanding ability of extended contexts while xRAG [7] integrates context compression by mapping documents into a compact token representation. Similarly, AAG [26] draws inspiration from human cognition, retrieving and recalling relevant knowledge to compensate for knowledge gaps. This approach activates relevant information within LLMs without relying on external resources. Our proposed method dynamically converts retrieved documents into parameters, effectively unifying internal and external knowledge while mitigating RAG hallucination issue. Furthermore, it functions as a plug-and-play solution to enhance the efficiency and adaptability of standard RAG systems by test-time parametric knowledge enhancement.
@@ 4.3 Main Results
In this section, we present the main experimental result and in-depth analysis of DyPRAG compared with vanilla model, standard RAG, efficient RAG, PRAG baselines and a combined setting . Notably, the vanilla model occasionally outperforms RAG in certain situations. We analyze the reasons for this in Appendix D, and confirm that it won’t affect the subsequent analysis.
@@ 4.4 Out-Of-Distribution (OOD) Performance
To further demonstrate the generalization ability of the DyPRAG method, we evaluate it in an out-of-distribution (OOD) scenario. Table 3 presents the OOD performance on the StrategyQA (SQA)[12] and IIRC[9] datasets. Compared to the standard RAG model, the vanilla model performs poorly due to the lack of relevant knowledge. In contrast, our DyPRAG successfully injects parametric knowledge, achieving comparable performance in SQA. Notably, DyPRAG-Combine which incorporates document-related parametric knowledge with ground-truth passages can answer these questions more correctly, achieving the best performance across all scenarios. For example, DyPRAG-Combine improves performance on SQA by 22.66% using LLaMA3.2-1B and on IIRC by 13.63% using LLaMA3-8B. These experimental results highlight the significant generalization ability of DyPRAG, surpassing that of PRAG, which is unable to handle this OOD scenario without additional offline training. Detailed implementation is provided in Appendix B.1 and analysis in Appendix D.
@@ 4.6 Analysis of Contextual and Parametric Knowledge Conflict and Fusion
When contextual knowledge and parametric knowledge conflict, LLMs may struggle to determine which source of information is more reliable [42; 51]. A key distinction between parametric RAG and standard RAG lies in whether external documents are used as contextual knowledge or transforming into parametric knowledge. Additionally, we aim to explore why combining parameter injection with in-context injection results in superior performance.
@@
In contrast, DyPRAG-Combine effectively integrates contextual knowledge with transformed parametric knowledge, enabling it to provide correct answers and demonstrating its ability to leverage both types of knowledge effectively. Compared to RAG, DyPRAG-Combine transforms the retrieved documents into parametric knowledge before injecting them into the model. This approach ensures that the LLM already contains relevant knowledge when answering questions, unlike RAG which often suffers from well-known hallucination issues [42; 40]. As a result, DyPRAG-Combine offers a practical and efficient solution to directly relieve the common hallucination problem in standard RAG, providing a plug-and-play improvement to enhance RAG’s performance in real-world applications.
@@
The retrieved documents in our experiments are primarily sourced from Wikipedia, which are already encountered by LLMs during the pre-training phase. In this section, we further investigate how DyPRAG performs on unseen contextual knowledge using the RAGTruth benchmark [33]. Specifically, we randomly sample 100 examples from the QA-type sub-dataset, which presents greater challenges (e.g., the required answers are only accessible in carefully crafted context). As shown in Figure 3, DyPRAG-Combine significantly outperforms RAG. This demonstrates that DyPRAG effectively enables LLMs to better internalize contextual knowledge and mitigate hallucinations, even when handling unseen data.  [^1] We provide further detail analysis of contextual and parametric knowledge in Appendix E.
@@ Appendix C Supplement Experiment Results
Can Perplexity Reflects Knowledge Conflict?
Recent studies have explored methods to detect hallucinations in LLMs and RAG systems by leveraging various metrics [6, 40]. Among these, we adopt the simplest yet effective metric, Perplexity (PPL) [35], to evaluate knowledge conflicts. As illustrated in Figure8, Vanilla and DyPRAG exhibit higher PPL, while DyPRAG-Combine and RAG demonstrate significantly lower PPL. However, these results are inconsistent with the findings in Table 1 and Table 3. For instance, although DyPRAG-Combine achieves the best performance on IIRC, its calculated PPL suggests a higher probability of knowledge conflicts, which is clearly incorrect.
We hypothesize that this discrepancy arises primarily from variations in model parameters caused by parameter injection in DyPRAG, which is undetectable with the current evaluation method. Developing more advanced methods to detect knowledge conflicts in DyPRAG remains an interesting avenue for future research.
@@
Appendix D Why Vanilla Outperforms RAG Occasionally ?
@@
In this section, we provide a detailed analysis of why the vanilla model occasionally outperforms RAG. As shown in Table 1, the vanilla model surpasses RAG most significantly in 2WQA, although the results vary across different models. For instance, the vanilla model outperforms RAG by 2.62 and 0.99 in Qwen2.5-1.5B and LLaMA3-8B on average, respectively. After analyzing the cases, we identify two key issues that affect RAG’s performance: 1) Poor Retriever. Following [39], we use BM25 as the retriever. However, in many cases, the retrieved documents contain only similar words rather than relevant content. This results in the provided content being unhelpful or even detrimental to the performance of LLMs. 2) Already Seen Data. During the pre-training stages of the selected LLMs [49, 31, 30], the external source we use (i.e., Wikipedia) has already been seen. This allows LLMs to answer certain questions independently, especially in simpler tasks like 2WQA. Moreover, the inclusion of incorrect or irrelevant context further degrades the performance of LLMs, as observed in Table 1.
@@
Appendix E Further Analysis of Contextual and Parametric Knowledge Conflict
@@
Appendix F Visualization of Parameter Translator Workflow.
@@
Appendix G Prompt for Main Experiments Evaluation
@@
Appendix H Prompt for Knowledge Internalization Evaluation
@@ 3. Relevance: Does it stay on topic while demonstrating depth?
Appendix I Limitations and Future Directions

### v3

@@ 1 Introduction
To mitigate incorrect internal knowledge and improve performance in QA, researchers have developed knowledge-augmented methods for LLMs. One of the most used methods called Retrieval-Augmented Generation (RAG) [13], which retrieves relevant documents from external sources (e.g., Wikipedia) and integrates both the retrieved documents and the query into the model’s input [17] (as shown in Figure 1 (a)). Standard RAG methods [3; 18; 39] share a common approach: injecting external knowledge into LLMs by appending retrieved documents to the input context, referred to as in-context injection. This technique harnesses the reasoning capabilities of LLMs to address knowledge gaps and reduce misinformation [5]. However, it introduces several challenges. As the number and length of retrieved documents increase, the extended context leads to significantly higher inference costs. Additionally, when the internal knowledge of LLMs conflicts with the external context, a phenomenon known as RAG hallucination (or knowledge conflict) arises, causing LLMs to produce incorrect answers even when accurate documents are provided [53; 42].
@@ 2.1 Retrieval Augmented Generation
Large language models (LLMs) have demonstrated remarkable performance across diverse applications. However, their inherent knowledge often falls short in handling knowledge-intensive tasks, highlighting the need for external knowledge integration to ensure robust performance in such contexts. A prominent approach to bridging this gap is Retrieval-Augmented Generation (RAG), which enhances LLMs by incorporating relevant external knowledge sources [4; 46; 47; 14]. The retrieved documents are appended to the LLM’s input context, enabling it to leverage knowledge beyond its training data [24]. We refer to this as the in-context injection paradigm. However, this approach leads to high inference costs and can exceed context length limits as the number and length of retrieved documents increase [50]. To address this issue, recent study introduces Parametric RAG (PRAG) [41], a paradigm that fine-tunes the model on augmented documents, encoding useful information into its parameters. While PRAG mitigates the inference cost, it introduces additional training and storage costs due to the need to obtain and store LoRA parameters. Our proposed method significantly reduces the high costs associated with standard RAG and PRAG while achieving superior generalization. By combining contextual knowledge with test-time generated parametric knowledge through DyPRAG, it enables LLMs to better manipulate knowledge, effectively mitigating RAG hallucination. This makes it a more practical and robust solution for real-world applications.
@@ 2.2 Context Compression
Context compression is widely adopted to improve the efficiency of LLMs in processing contextual knowledge. Recent studies [34; 11; 8] propose condensing long contexts into soft prompts, allowing LLMs to utilize information more effectively. Meanwhile, other works  [31; 48] focus on transforming context chunk into LoRA modules to improve the understanding ability of extended contexts while xRAG [7] integrates context compression by mapping documents into a compact token representation. Similarly, AAG [26] draws inspiration from human cognition, retrieving and recalling relevant knowledge to compensate for knowledge gaps. This approach activates relevant information within LLMs without relying on external resources. Our proposed method dynamically converts retrieved documents into parameters, effectively unifying internal and external knowledge while mitigating RAG hallucination issue. Furthermore, it functions as a plug-and-play solution to enhance the efficiency and adaptability of standard RAG systems by test-time parametric knowledge enhancement.
@@ 4.3 Main Results
In this section, we present the main experimental result and in-depth analysis of DyPRAG compared with vanilla model, standard RAG, efficient RAG, PRAG baselines and a combined setting . Notably, the vanilla model occasionally outperforms RAG in certain situations. We analyze the reasons for this in Appendix E, and confirm that it won’t affect the subsequent analysis.
@@ 4.4 Out-Of-Distribution (OOD) Performance
To further demonstrate the generalization ability of the DyPRAG method, we evaluate it in an out-of-distribution (OOD) scenario. Table 3 presents the OOD performance on the StrategyQA (SQA)[12] and IIRC[9] datasets. Compared to the standard RAG model, the vanilla model performs poorly due to the lack of relevant knowledge. In contrast, our DyPRAG successfully injects parametric knowledge, achieving comparable performance in SQA. Notably, DyPRAG-Combine which incorporates document-related parametric knowledge with ground-truth passages can answer these questions more correctly, achieving the best performance across all scenarios. For example, DyPRAG-Combine improves performance on SQA by 22.66% using LLaMA3.2-1B and on IIRC by 13.63% using LLaMA3-8B. These experimental results highlight the significant generalization ability of DyPRAG, surpassing that of PRAG, which is unable to handle this OOD scenario without additional offline training. Detailed implementation is provided in Appendix B.1 and analysis in Appendix E.
@@ 4.6 Analysis of Contextual and Parametric Knowledge Conflict and Fusion
When contextual knowledge and parametric knowledge conflict, LLMs may struggle to determine which source of information is more reliable [44; 53]. A key distinction between parametric RAG and standard RAG lies in whether external documents are used as contextual knowledge or transforming into parametric knowledge. Additionally, we aim to explore why combining parameter injection with in-context injection results in superior performance.
@@
In contrast, DyPRAG-Combine effectively integrates contextual knowledge with transformed parametric knowledge, enabling it to provide correct answers and demonstrating its ability to leverage both types of knowledge effectively. Compared to RAG, DyPRAG-Combine transforms the retrieved documents into parametric knowledge before injecting them into the model. This approach ensures that the LLM already contains relevant knowledge when answering questions, unlike RAG which often suffers from well-known hallucination issues [44; 42]. As a result, DyPRAG-Combine offers a practical and efficient solution to directly relieve the common hallucination problem in standard RAG, providing a plug-and-play improvement to enhance RAG’s performance in real-world applications.
@@
The retrieved documents in our experiments are primarily sourced from Wikipedia, which are already encountered by LLMs during the pre-training phase. In this section, we further investigate how DyPRAG performs on unseen contextual knowledge using the RAGTruth benchmark [35]. Specifically, we randomly sample 100 examples from the QA-type sub-dataset, which presents greater challenges (e.g., the required answers are only accessible in carefully crafted context). As shown in Figure 3, DyPRAG-Combine significantly outperforms RAG. This demonstrates that DyPRAG effectively enables LLMs to better internalize contextual knowledge and mitigate hallucinations, even when handling unseen data.  [^1] We explore metrics for detecting hallucinations in Retrieval-Augmented Generation (RAG) in Appendix LABEL:app:rag_hal. Additionally, we present a detailed analysis of contextual and parametric knowledge in Appendix F.
@@ Appendix C Supplement Experiment Results
Appendix D Exploring Metrics for RAG Hallucination Detection
Can Perplexity Reflects Knowledge Conflict?
Recent studies have explored methods to detect hallucinations in LLMs and RAG systems by leveraging various metrics [6, 42]. Among these, we first adopt the simplest yet effective metric which only need single generation, Perplexity (PPL) [37], to evaluate knowledge conflicts. As illustrated in Figure8, Vanilla and DyPRAG exhibit higher PPL, while DyPRAG-Combine and RAG demonstrate significantly lower PPL. However, these results are inconsistent with the findings in Table 1 and Table 3. For instance, although DyPRAG-Combine achieves the best performance on IIRC, its calculated PPL suggests a higher probability of knowledge conflicts, which is clearly incorrect. We hypothesize that this discrepancy primarily stems from variations in model parameters introduced by parameter injection in DyPRAG, which cannot be detected using the simple PPL method. Given that different tokens contribute unequally to the overall semantics of a sentence, the PPL, which calculated as the average of token-level uncertainty, fails to effectively capture the uncertainty of the entire sequence.
Effective Detection with Sentence-Level Metrics.
Given the limitations of PPL, we decided to explore alternative metrics that leverage multiple generations. Research has shown that generating multiple outputs for a single input is beneficial for estimating sequence-level uncertainty. To this end, we set the temperature to 1.0, top_p to 0.95, and top_k to 20, generating five responses to calculate Entropy (EN), Length Normalized Entropy (LEN) [29], and Lexical Similarity (LS) [27] to evaluate the probability of RAG hallucination [^2] We use the implementation in https://github.com/alibaba/eigenscore. As shown in Table 8, our approach demonstrates reduced knowledge conflict in most scenarios, especially in our strongest DyPRAG-Combine.
We observed that both EN and LEN increase when in-context injection is applied, suggesting that in RAG systems, the retrieved passages often conflict with the model’s internal knowledge. In contrast, utilizing DyPRAG to inject converted parametric knowledge significantly reduces the likelihood of knowledge conflict, demonstrating the effectiveness of DyPRAG. However, the LS results indicate that adding context reduces conflict, which contradicts the established definition of RAG hallucination. We argue that EN and LEN are more suitable for effective RAG hallucination detection in DyPRAG settings. Exploring more effective detection methods remains an important direction for future work.
Table 8: We present the experimental results for the knowledge conflict metrics of DyPRAG and DyPRAG-Combine, in comparison with Vanilla and Standard RAG. Three multi-generation metrics are employed: Entropy (EN), Length Normalized Entropy (LEN), and Lexical Similarity (LS). In these metrics, $\uparrow$ indicates that higher values are better, while $\downarrow$ indicates the opposite. The best performance for each metric is highlighted in bold. The backbone model is the LLaMA3.2-1B.
Metric Method 2WQA (total) HQA (total) PQA CWQ SQA IIRC
EN $\downarrow$ Vanilla  3.187  3.176  3.251  3.163  3.178  3.011
DyPRAG (ours)  2.199  1.999  1.757  2.860  2.805  2.544
[…]
DyPRAG-Combine (ours)  2.755  2.470  3.584  3.467  3.136  2.555
LEN $\downarrow$ Vanilla  0.637  0.635  0.650  0.633  0.636  0.602
DyPRAG (ours)  0.440  0.400  0.586  0.572  0.561  0.509
[…]
DyPRAG-Combine (ours)  0.551  0.494  0.719  0.693 0.627 0.511
LS $\uparrow$ Vanilla  0.923  0.936  0.723  0.730  0.497  0.963
DyPRAG (ours)  0.915  0.933  0.842  0.859  0.527  0.966
[…]
DyPRAG-Combine (ours)  0.953  0.959  0.966  0.988  0.853  0.975
Appendix E Why Vanilla Outperforms RAG Occasionally ?
@@
In this section, we provide a detailed analysis of why the vanilla model occasionally outperforms RAG. As shown in Table 1, the vanilla model surpasses RAG most significantly in 2WQA, although the results vary across different models. For instance, the vanilla model outperforms RAG by 2.62 and 0.99 in Qwen2.5-1.5B and LLaMA3-8B on average, respectively. After analyzing the cases, we identify two key issues that affect RAG’s performance: 1) Poor Retriever. Following [41], we use BM25 as the retriever. However, in many cases, the retrieved documents contain only similar words rather than relevant content. This results in the provided content being unhelpful or even detrimental to the performance of LLMs. 2) Already Seen Data. During the pre-training stages of the selected LLMs [51, 33, 32], the external source we use (i.e., Wikipedia) has already been seen. This allows LLMs to answer certain questions independently, especially in simpler tasks like 2WQA. Moreover, the inclusion of incorrect or irrelevant context further degrades the performance, as observed in Table 1.
@@
Appendix F Further Analysis of Contextual and Parametric Knowledge Conflict
@@
Appendix G Visualization of Parameter Translator Workflow.
@@
Appendix H Prompt for Main Experiments Evaluation
@@
Appendix I Prompt for Knowledge Internalization Evaluation
@@ 3. Relevance: Does it stay on topic while demonstrating depth?
Appendix J Limitations and Future Directions